OpenStudy (anonymous):

The time required to finish a test in normally distributed with a mean of 80 minutes and a standard deviation of 15 minutes. What is the probability that a student chosen at random will finish the test in more than 110 minutes?

OpenStudy (anonymous):

\[P(X>110)=P\left(\frac{X-80}{15}>\frac{110-80}{15}\right)=P(Z>2)\]

OpenStudy (anonymous):

thank you.!

OpenStudy (anonymous):

yw